OpenStudy (anonymous):

you invest 1,000 at 5% interest componded annually. in how many years will you have 2000 in your account

OpenStudy (anonymous):

maybe this can help http://www.moneychimp.com/articles/finworks/fmfutval.htm

OpenStudy (anonymous):

ok so we have p = 1000, r = 5, t = ?, I = 2000

OpenStudy (anonymous):

I = p (1 + r/100)^n

OpenStudy (anonymous):

2000 = 1000(1.05)^n

OpenStudy (anonymous):

(1.05)^n = 2

OpenStudy (anonymous):

ok i am wrong somewhere

OpenStudy (anonymous):

thank you.